Two point charges of 2 micro C and + 6 micro C repel each other with a force of 12 newton if each given an additional charge of -4 micro C then force will become
 

Grade:12th pass

2 Answers

Ankita
4 years ago
Let q=2 micro C
      Q= 6 micro C
Given, F = 12 N
kqQ/r2 = 12
r= 9×109×2×10-6×6×10-6 /12
r2 = 9×10-3 m
 
Q'= 6-4= 2 micro C 
q'= 2-4 = -2 micro C
 
F= 9×109× 2×10-6×2×10-6 / 9×10-3
F = 4 N
Khimraj
4 years ago
 
kqQ/r2 = 12
r= 9×109×2×10-6×6×10-6 /12
r2 = 9×10-3 m
F= 9×109× 2×10-6×2×10-6 / 9×10-3
F = 4 N
…...................................